# Information Security Based on LLM Approaches: A Review

### v1

@@ 3. Malware Detection
Malware behavior identification aims to detect potential threats by analyzing the operation patterns of programs (e.g., file access, registry modification, network communication, and process creation, etc.). In recent years, large language models (LLMs) have been gradually introduced into malware detection tasks due to their excellent semantic understanding and modeling capabilities. Compared with traditional machine learning methods that rely on a large number of labeled samples, feature engineering, and computational resources, detection methods based on large language models are able to learn complex behavioral semantic patterns directly from raw data and identify unknown malicious behaviors more efficiently. By fusing pre-trained language models with behavioral sequence modeling techniques, the method shows significant advantages in improving detection accuracy and reducing false alarm rate and becomes an important means to build an intelligent malware detection system(liu2024gastrace, ).
In recent years, researchers have been exploring the application of large language models and deep learning methods to malware detection and classification tasks, and have achieved remarkable results. Qiao(qiao2021malware, ) et al. propose a malware classification method that combines Word2Vec and multilayer perceptron (MLP), and the model performs well in both traditional and IoT malware classification tasks by converting byte sequences into vectors and inputting them into a deep neural network. Victor et al. experimentally verified the potential of ChatGPT for malware identification by driving a scanner that successfully detects multiple malicious processes and avoids interference from benign processes(kasneci2023chatgpt, ). Alzaidy et al. applied RNNs and CNNs to classify Windows executable files and analyzed the impact of adversarial samples based on the results. Alzaidy et al. apply RNN and CNN to Windows executable file classification and analyze the impact of adversarial examples based on the results, which show that JSMA and C&W attacks can circumvent the model detection to a certain extent, and the robustness of the attacks varies among different models(alzaidy2024adversarial, ). Marwaha et al. propose converting the binary files of Android malware into RGB images for visual classification using the VGG16 network. This method significantly improves classification accuracy and the F1 score, highlighting the unique advantages of visual features in malware detection(marwaha2023retracted, ). The binary to malware image flow is shown in Fig. 5.
@@ 4.1. DDoS Attack Prediction
With its powerful feature learning and generalization capabilities, LLM provides technical support for the efficient prediction of DDoS attacks. Through in-depth modeling of massive network traffic data, large models can automatically identify potential patterns between normal and abnormal traffic, thus upgrading the security strategy from a traditional responsive mechanism to an active prediction and defense mechanism(wu2025exploring, ). Feature extraction is the core aspect of the large model that plays a role in DDoS attack prediction. Using deep neural networks such as Convolutional Neural Networks (CNN), Recurrent Neural Networks (RNN), and Long Short-Term Memory Networks (LSTM), key features reflecting the state of the network, such as packet sizes, transmission intervals, communication protocols, etc., can be automatically learned from the original network stream. These models are particularly good at handling sequence data with time-dependent relationships, and thus exhibit high accuracy and robustness in anomaly detection(li2024detecting, ).
@@
In practical research, Cheon et al proposed a deep learning method based on Word2Vec for DDoS attack detection(cheon2022novel, ). The method significantly improves the model’s ability to capture data patterns by encoding network traffic into semantic embeddings, which effectively enhances the detection performance. Meanwhile, Guastalla et al proposed an LLM detection strategy that combines Few-shot cue learning, Fine-tuning of the model, and neural network architecture. This strategy generates semantic understanding and fine-tuning optimization through LLM under the premise of using a small number of labeled samples, and ultimately achieves higher detection accuracy than traditional neural networks, which provides an innovative direction for DDoS defense in resource-constrained scenarios(guastalla2023application, ). The results of DDoS attack detection in practice are as follows: Listing 1 and Tab 1.
@@ 4.2. Network Security Log Analysis
It has been shown that deep learning and natural language processing technologies have good performance in log anomaly detection. The LogFiT model proposed by Crispin et al. is based on BERT for pre-training and fine-tuning to learn the linguistic structure of system logs(almodovar2022can, ). The model optimizes the normal logging pattern through two objectives: mask prediction and center-of-mass distance minimization, and combines top-k prediction accuracy and center-of-mass distance in the inference stage to determine whether the log is abnormal or not. Experimental results show that LogFiT exhibits good accuracy and robustness in the log anomaly detection task.
@@ 4.3. Network Abnormal Traffic Detection
LLM is able to construct language models for network packets and learn the feature distribution and behavioral patterns of normal traffic. When abnormal traffic occurs, patterns that deviate from normal behavior can be identified through model comparison, achieving efficient and accurate abnormality detection. In addition, LLM can also mine the deep dependencies in the traffic to assist analysts in more accurate threat identification. The FlowTransformer model proposed by Manocchio et al(manocchio2024flowtransformer, ) represents network traffic as a sequence of feature vectors (including source/destination IPs, protocol types, etc.) and uses the Transformer encoder to model its temporal dependencies to achieve classification of normal and abnormal traffic. Experiments show that this method has high accuracy and generalization ability in anomaly detection tasks. In addition, Chai et al. proposed a detection method combining CNN and Transformer, which first extracts local features by a convolutional neural network, then models global dependencies by a Transformer encoder, and finally expresses them as abstract features by a decoder, and completes the classification with the help of SVM. The method performs well in recognizing unknown attacks and potential threats, and has good detection accuracy and practical value (chai2023ctsf, ).
@@ 5. Security Vulnerability Detection
Security vulnerability detection relies on an in-depth understanding of program semantics and structure, so it is crucial to select appropriate language models. Currently, commonly used models include n-gram, RNN, LSTM, and models based on the Transformer architecture, such as BERT and GPT. The Transformer model has become a mainstream solution due to its powerful semantic modeling capability. In practice, researchers usually use large-scale pre-trained language models and fine-tune them with data from the security domain to improve the performance of the models in vulnerability detection tasks. The training process often employs a transfer learning strategy to migrate the linguistic knowledge of the general-purpose model to a specific security corpus, so as to improve the model’s ability to generalize the vulnerability features.
For example, the CodeSentry framework proposed by Jones et al(jones2024codesentry, ) is based on GPT-2, which encodes words and sub-words of C, C++, and Java code, and identifies the vulnerability types through the output vectors, which significantly reduces manual feature engineering.SmartConDetect, designed by Jeon et al., utilizes the BERT model to semantically analyze the Solidity smart contract. SmartConDetect by Jeon et al(jeon2021smartcondetect, ). utilizes the BERT model to semantically analyze Solidity smart contracts and realizes automated vulnerability detection through softmax classifiers, while Szabó et al. implement static analysis in combination with the GPT API, and detect sensitive snippets of front-end applications through natural language hints to effectively identify vulnerabilities such as CWE-653(szabo2023new, ).
@@ 6.1. Malicious Code Detection
The introduction of large language models (LLMs) provides a breakthrough in malicious code analysis. With their powerful capabilities in semantic modeling, context understanding, and sequence learning, LLMs can extract deeper semantic features from complex instruction sequences and API call paths, making up for the shortcomings of traditional feature engineering(li2024cobra, ). At the same time, the large model can generalize the identification of variant codes, which can effectively explore the latent malicious behavior patterns and significantly improve the accuracy and adaptability of detection.
@@
In the classification and detection stage, classification algorithms incorporating features extracted from LLMs have demonstrated superior performance. Combined with deep learning or supervised learning methods, researchers can build high-precision classifiers based on categories such as viruses, worms, Trojans, etc., to quickly recognize unknown malicious samples. In addition, LLMs can also take on the core modeling task in the end-to-end detection framework, which improves the robustness and practicality of the system in real attack and defense environments(patsakis2024assessing, ). The Fig. 8 shows the core process of malicious code detection and highlights the key role of Large Language Models (LLMs) in feature extraction and semantic understanding. By fusing static and dynamic features, large models are able to model complex behavioral logic and improve detection accuracy.
@@ 6.2. Cryptographic Algorithms
With the continuous evolution of artificial intelligence technology, the application of large language models (LLMs) in symmetric encryption and asymmetric encryption algorithms has become a new hot spot in network security research. In symmetric encryption algorithms, LLMs can play a role in the following key aspects: first, the optimization of the key exchange mechanism, using its generation and prediction capabilities to assist in the establishment of a more secure key negotiation process; second, the improvement of encryption and decryption efficiency, through the identification of data patterns and contextual features, to improve the processing efficiency of the algorithm; third, the identification of vulnerabilities and risk assessment, the model can be used to analyze the history of the attack samples and the description of the protocol, to assist in discovering potential security risks. The model can analyze historical attack samples and protocol descriptions to assist in the discovery of potential security risks(gupta2024enhancing, ).

### v2

@@ 3. Malware Detection
Malware behavior identification aims to detect potential threats by analyzing the operation patterns of programs (e.g., file access, registry modification, network communication, and process creation, etc.). In recent years, large language models (LLMs) have been gradually introduced into malware detection tasks due to their excellent semantic understanding and modeling capabilities(Bensaoud et al., 2024). Compared with traditional machine learning methods that rely on a large number of labeled samples, feature engineering, and computational resources, detection methods based on large language models are able to learn complex behavioral semantic patterns directly from raw data and identify unknown malicious behaviors more efficiently. By fusing pre-trained language models with behavioral sequence modeling techniques, the method shows significant advantages in improving detection accuracy and reducing false alarm rate and becomes an important means to build an intelligent malware detection system(Liu et al., 2024).
In recent years, researchers have been exploring the application of large language models and deep learning methods to malware detection and classification tasks, and have achieved remarkable results. Qiao(Qiao et al., 2021) et al. propose a malware classification method that combines Word2Vec and multilayer perceptron (MLP), and the model performs well in both traditional and IoT malware classification tasks by converting byte sequences into vectors and inputting them into a deep neural network. Victor et al. experimentally verified the potential of ChatGPT for malware identification by driving a scanner that successfully detects multiple malicious processes and avoids interference from benign processes(Kasneci et al., 2023). Alzaidy et al. applied RNNs and CNNs to classify Windows executable files and analyzed the impact of adversarial samples based on the results. Alzaidy et al. apply RNN and CNN to Windows executable file classification and analyze the impact of adversarial examples based on the results, which show that JSMA and C&W attacks can circumvent the model detection to a certain extent, and the robustness of the attacks varies among different models(Alzaidy and Binsalleeh, 2024). Marwaha et al. propose converting the binary files of Android malware into RGB images for visual classification using the VGG16 network. This method significantly improves classification accuracy and the F1 score, highlighting the unique advantages of visual features in malware detection(Marwaha et al., 2023). The binary to malware image flow is shown in Fig. 5.
@@ 4.1. DDoS Attack Prediction
With its powerful feature learning and generalization capabilities, LLM provides technical support for the efficient prediction of DDoS attacks. Through in-depth modeling of massive network traffic data, large models can automatically identify potential patterns between normal and abnormal traffic, thus upgrading the security strategy from a traditional responsive mechanism to an active prediction and defense mechanism(Wu et al., 2025). Feature extraction is the core aspect of the large model that plays a role in DDoS attack prediction. Using deep neural networks such as Convolutional Neural Networks (CNN), Recurrent Neural Networks (RNN), and Long Short-Term Memory Networks (LSTM), key features reflecting the state of the network, such as packet sizes, transmission intervals, communication protocols, etc., can be automatically learned from the original network stream. These models are particularly good at handling sequence data with time-dependent relationships, and thus exhibit high accuracy and robustness in anomaly detection(Li et al., 2024d).
@@
In practical research, Cheon et al proposed a deep learning method based on Word2Vec for DDoS attack detection(Cheon et al., 2022). The method significantly improves the model’s ability to capture data patterns by encoding network traffic into semantic embeddings, which effectively enhances the detection performance. Meanwhile, Guastalla et al proposed an LLM detection strategy that combines Few-shot cue learning, Fine-tuning of the model, and neural network architecture. This strategy generates semantic understanding and fine-tuning optimization through LLM under the premise of using a small number of labeled samples, and ultimately achieves higher detection accuracy than traditional neural networks, which provides an innovative direction for DDoS defense in resource-constrained scenarios(Guastalla et al., 2023). The results of DDoS attack detection in practice are as follows: Listing 1 and Tab 1.
@@ 4.2. Network Security Log Analysis
It has been shown that deep learning and natural language processing technologies have good performance in log anomaly detection. The LogFiT model proposed by Crispin et al. is based on BERT for pre-training and fine-tuning to learn the linguistic structure of system logs(Almodovar et al., 2022). The model optimizes the normal logging pattern through two objectives: mask prediction and center-of-mass distance minimization, and combines top-k prediction accuracy and center-of-mass distance in the inference stage to determine whether the log is abnormal or not. Experimental results show that LogFiT exhibits good accuracy and robustness in the log anomaly detection task.
@@ 4.3. Network Abnormal Traffic Detection
LLM is able to construct language models for network packets and learn the feature distribution and behavioral patterns of normal traffic. When abnormal traffic occurs, patterns that deviate from normal behavior can be identified through model comparison, achieving efficient and accurate abnormality detection. In addition, LLM can also mine the deep dependencies in the traffic to assist analysts in more accurate threat identification. The FlowTransformer model proposed by Manocchio et al(Manocchio et al., 2024) represents network traffic as a sequence of feature vectors (including source/destination IPs, protocol types, etc.) and uses the Transformer encoder to model its temporal dependencies to achieve classification of normal and abnormal traffic. Experiments show that this method has high accuracy and generalization ability in anomaly detection tasks. In addition, Chai et al. proposed a detection method combining CNN and Transformer, which first extracts local features by a convolutional neural network, then models global dependencies by a Transformer encoder, and finally expresses them as abstract features by a decoder, and completes the classification with the help of SVM. The method performs well in recognizing unknown attacks and potential threats, and has good detection accuracy and practical value (Chai et al., 2023).
@@ 5. Security Vulnerability Detection
Security vulnerability detection relies on an in-depth understanding of program semantics and structure, so it is crucial to select appropriate language models. Currently, commonly used models include n-gram, RNN, LSTM, and models based on the Transformer architecture, such as BERT and GPT. The Transformer model has become a mainstream solution due to its powerful semantic modeling capability. In practice, researchers usually use large-scale pre-trained language models and fine-tune them with data from the security domain to improve the performance of the models in vulnerability detection tasks. The training process often employs a transfer learning strategy to migrate the linguistic knowledge of the general-purpose model to a specific security corpus, so as to improve the model’s ability to generalize the vulnerability features(Bennouk et al., 2024).
For example, the CodeSentry framework proposed by Jones et al(Jones and Omar, 2024) is based on GPT-2, which encodes words and sub-words of C, C++, and Java code, and identifies the vulnerability types through the output vectors, which significantly reduces manual feature engineering.SmartConDetect, designed by Jeon et al., utilizes the BERT model to semantically analyze the Solidity smart contract. SmartConDetect by Jeon et al(Jeon et al., 2021). utilizes the BERT model to semantically analyze Solidity smart contracts and realizes automated vulnerability detection through softmax classifiers, while Szabó et al. implement static analysis in combination with the GPT API, and detect sensitive snippets of front-end applications through natural language hints to effectively identify vulnerabilities such as CWE-653(Szabó and Bilicki, 2023).
@@ 6.1. Malicious Code Detection
The introduction of large language models (LLMs) provides a breakthrough in malicious code analysis. With their powerful capabilities in semantic modeling, context understanding, and sequence learning, LLMs can extract deeper semantic features from complex instruction sequences and API call paths, making up for the shortcomings of traditional feature engineering(Li et al., 2024b). At the same time, the large model can generalize the identification of variant codes, which can effectively explore the latent malicious behavior patterns and significantly improve the accuracy and adaptability of detection.
@@
In the classification and detection stage, classification algorithms incorporating features extracted from LLMs have demonstrated superior performance. Combined with deep learning or supervised learning methods, researchers can build high-precision classifiers based on categories such as viruses, worms, Trojans, etc., to quickly recognize unknown malicious samples. In addition, LLMs can also take on the core modeling task in the end-to-end detection framework, which improves the robustness and practicality of the system in real attack and defense environments(Patsakis et al., 2024). The Fig. 8 shows the core process of malicious code detection and highlights the key role of Large Language Models (LLMs) in feature extraction and semantic understanding. By fusing static and dynamic features, large models are able to model complex behavioral logic and improve detection accuracy.
@@ 6.2. Cryptographic Algorithms
With the continuous evolution of artificial intelligence technology, the application of large language models (LLMs) in symmetric encryption and asymmetric encryption algorithms has become a new hot spot in network security research. In symmetric encryption algorithms, LLMs can play a role in the following key aspects: first, the optimization of the key exchange mechanism, using its generation and prediction capabilities to assist in the establishment of a more secure key negotiation process; second, the improvement of encryption and decryption efficiency, through the identification of data patterns and contextual features, to improve the processing efficiency of the algorithm; third, the identification of vulnerabilities and risk assessment, the model can be used to analyze the history of the attack samples and the description of the protocol, to assist in discovering potential security risks. The model can analyze historical attack samples and protocol descriptions to assist in the discovery of potential security risks(Gupta et al., 2024).